OpenStudy (anonymous):

Which of the following statements have the same results?: I: f(1) when f(x)=5x+1 II: f^-1(3) when f(x)=2x+3/5 III: 3y-7=y+5

OpenStudy (ksaimouli):

substitute each value in x for example= f(1)= 5(1)+1= 6

OpenStudy (tkhunny):

Is the second one \(2x + \dfrac{3}{5}\), as you have written it, or did you intend \(\dfrac{2x+3}{5}\)?

OpenStudy (anonymous):

For the first one I got 6. But I’m not sure with the others… This is the second one: 

OpenStudy (tkhunny):

Do you see how that is NOT what you originally wrote?

OpenStudy (anonymous):

Sorry, it's what I meant to write.

OpenStudy (tkhunny):

You didn't answer my question. It is important to know how and why it is different. f(1) when f(x)=5x+1 ==> 5(1) + 1 = 5 + 1 = 6 -- We've seen that, already.

OpenStudy (tkhunny):

3y - 7 = y + 5 Subtract y 2y - 7 = 5 Add 7 2y = 12 Divide by 2 y = 6

OpenStudy (tkhunny):

The parentheses I added are NOT optional. Remember your Order of Operations. f^-1(3) when f(x)=(2x+3)/5 3 = (2x+3)/5 Multiply by 5 15 = 2x + 3 Subtract 3 12 = 2x Divide by 2 6 = x They all seem to be coming out 6.

OpenStudy (anonymous):

I see the difference now, I'll keep that in mind. Thanks for your help:)